angle A= 47.3, side b= 14.0, and c=36.Find,angle C.... i got 33.7 degrees and pretty sure that's wrong!! please work it out for me!
Answer by stanbon(75887)

angle A= 47.3, side b= 14.0, and c=36.Find,angle C.... i got 33.7 degrees and pretty sure that's wrong!
----
Law of Cosines::
a^2 = b^2 + c^2 - 2bc*cos(A)
----
a^2 = 14^2 + 36^2 - 2*14*36*cos(47.3 deg) = 808.42
a = 28.43
---------------
Law of Sines::
sin(C)/36 = sin(47.3)/28.43
-----
sin(C) = [36/28.43]*sin(47.3) = 0.9305
---
Ans: C = arcsin(0.9305) = 68.51 deg
-----------------
Cheers,
Stan H.
===============